# GSON: A Group-based Social Navigation Framework with Large Multimodal Model

### v2

@@ Abstract
With the increasing presence of service robots and autonomous vehicles in human environments, navigation systems need to evolve beyond simple destination reach to incorporate social awareness. This paper introduces GSON, a novel group-based social navigation framework that leverages Large Multimodal Models (LMMs) to enhance robots’ social perception capabilities. Our approach uses visual prompting to enable zero-shot extraction of social relationships among pedestrians and integrates these results with robust pedestrian detection and tracking pipelines to overcome the inherent inference speed limitations of LMMs. The planning system incorporates a mid-level planner that sits between global path planning and local motion planning, effectively preserving both global context and reactive responsiveness while avoiding disruption of the predicted social group. We validate GSON through extensive real-world mobile robot navigation experiments involving complex social scenarios such as queuing, conversations, and photo sessions. Comparative results show that our system significantly outperforms existing navigation approaches in minimizing social perturbations while maintaining comparable performance on traditional navigation metrics. Video demonstration: https://youtu.be/mqkMAY0yAAk.
@@ I Introduction
The growth of service robots has driven significant research on autonomous systems capable of navigating human-centered environments [1, 2, 3]. However, a critical gap exists in current navigation systems: while they excel at trajectory prediction and obstacle avoidance [4, 5, 6, 7, 8], they often fail to recognize and respect complex social contexts within crowds, such as photography sessions or queuing behaviors, as illustrated in Fig. 1. In the broader context of social robot navigation [9, 10], the goal is not only for the robot to reach its destination, but also to interact appropriately with humans without degrading their experience. In this work, we focus on developing a robot navigation system that can understand and respect the contextual social interactions, where we represent as social groups. These social groups represent meaningful patterns of semantic interaction among individuals within the crowd.
Figure 1: The robot should recognize the social context and navigate around the queue instead of walking through it.
The core issue in building such a socially aware navigation system is how to accurately identify contextual social interactions in dynamic and unpredictable crowds, and exploit these social interactions to guide the planning system. For the perception, accurately identifying contextual social interactions in dynamic, unpredictable crowds requires semantic understanding beyond simple proximity or velocity patterns. Traditional methods based on predefined rules or in-domain learning [4, 5, 11, 12, 13, 14, 15, 16] struggle in open-world environments due to data scarcity and high visual variability. Regarding planning, generating socially appropriate motions that respect these interactions demands planning at the right abstraction level—traditional local planners often fall into local minima due to short planning horizons [17, 18, 19, 20], while global path planners cannot adapt quickly to dynamic social interactions. In contrast to modular systems, end-to-end solutions aim to bypass the traditional perception-planning decomposition [21, 22, 23, 24]. However, these approaches often overfit demonstrations and struggle to interpret complex contextual information. More recently, some works have explored the use of large multimodal models (LMMs) to select sampled trajectories [25] or high-level directions[26, 27, 28] to improve social awareness without explicit social interactions modeling. However, these methods are inefficient as they require querying the LMM at every planning step, and the resulting social behaviors are often not interpretable.
@@ II Related Work
Significant research has focused on social group detection in crowded pedestrian environments, which can be approached from both exo-centric and ego-centric perspectives. The exo-centric approach [29] employs external sensors to provide a comprehensive top-down view for crowd analysis. Various methods have been proposed for exo-centric group detection to support motion tracking [30, 31], trajectory prediction[12, 32, 33], group activity recognition[34, 35, 36], and crowd analysis [37, 38, 39]. However, this approach has two key limitations: it relies on external sensors and offline processing, and it often overlooks the rich semantics of crowd interactions by focusing mainly on motion patterns or domain-specific overfitting. For ego-centric group detection, prior approaches to crowd grouping have employed clustering based on motion patterns [40, 41, 13, 14] and distance [42, 43, 14], or learning group information in a specific dataset using graph neural network [15, 16], predicting cost function[44], or leveraging F-formation theories[45]. While these methods excel at robot-centric perception, they still face limitations in capturing semantic context and tend to overfit specific scenarios. In GSON, we leverage the zero-shot visual reasoning capabilities of the Large Multimodal Model (LMM) to enable semantic reasoning of social groups given ego-centric observation.
@@ II-B Social Robot Navigation
Previous studies have proposed various methods to improve collision avoidance in robot navigation [18, 5, 6, 8]. With the increasing number of mobile robots in human environments, there’s a need beyond collision avoidance to improve social awareness. Traditionally, researchers have applied physics-inspired approaches such as the Social Force Model [46] to robot navigation [4, 47, 48]. However, SFM-based methods only implicitly represent social awareness as pedestrian motion and suffer from limited generalization due to parameter selection. Group-based navigation strategies integrate dynamic human grouping into social navigation. Methods that combine group information with model-based planners [40, 43, 14] or reinforcement learning controllers [49, 16] show that representing human groups improves robot safety and social awareness. However, most group-based methods define groups only in terms of position and velocity, ignoring the semantic social interaction between pedestrians. Another line of research uses imitation learning [21, 22, 23, 24] to implicitly learn social awareness from demonstrations. However, due to the data scarcity problem, the learned policies capture only limited behaviors in the dataset. Recent foundation model advances have integrated large multimodal models for generating high-level actions [27, 28], scoring on trajectories [25] or directions [26], and distilling cost into imitation learning [50]. These methods are inefficient, requiring LMM queries at every planning step, with often uninterpretable social behaviors. Our approach uses LMMs only for low-frequency semantic social group detection when new pedestrians appear in observation, handling high-frequency tracking and planning without querying large models, creating a more practical system.
@@ III Problem Formulation
We consider the problem of social robot navigation in a 2D Euclidean space $\mathbb{R}^{2}$. This space is populated by $n$ dynamic pedestrians $H_{i}$ with a specific social structure. We consider the social structure as $m$ social groups $G_{j}$. The robot is initialized with a 2D occupancy map $M$ at the start position $S_{s}$ and a target position $S_{g}$ in the map. Given the 2D lidar point cloud $P^{t}$ and the RGB image observation $I^{t}$, the goal is to enable the robot to safely navigate to the goal position while minimizing the time it spends trespassing any social group. The robot must infer the social groups of pedestrians from the limited observation and generate an appropriate trajectory $T$. We focus on designing the method to estimate the social group with the aid of the Large Multimodal Model and integrate the estimated result into the navigation system.
@@ IV-B Social Group Estimation with Large Multimodal Models
The core idea of GSON is to model contextual social interactions in dynamic and unpredictable crowds using social groups, which then guide the planning system. A social group in this context refers to two or more individuals who are interacting, conversing, or moving together with shared intentions or social connections. To identify these social groups, we first build a robust pipeline for pedestrian detection and tracking. After obtaining detection and tracking results of pedestrians, identifying social groups requires commonsense knowledge about social interaction. Therefore, we leverage the multi-modal reasoning capabilities of the LMM and employ visual prompting to identify social groups.
@@
where $\mathcal{H}_{j}^{proj}=\big{\{}p_{i}^{3D}\,\big{|}\,\pi(p_{i}^{3D})\in B_{j}\big{\}}$ is the set of point cloud coordinates whose projections fall within the bounding box $B_{j}$, $\pi:\mathbb{R}^{3}\to\mathbb{R}^{2}$ denotes the projection from 3D point cloud coordinates to 2D image coordinates,$\mathbf{c}_{j}=(c_{j}^{x},c_{j}^{y})$ is the center coordinates of bounding box $B_{j}$, $\|\cdot\|_{2}$ represents the Euclidean distance.
After this filtering process, we obtain refined pedestrian positions in 2D coordinates $p_{i}=[x_{i},y_{i}]$, where the $z$ dimension is omitted as we focus on the ground plane for navigation purposes. Finally, we apply the Hungarian algorithm for data association and Kalman filtering for estimating and predicting the position of each pedestrian with a unique tracking ID, enabling robust tracking for subsequent social group detection. In this way, we detect and track the bounding box and tracking ID of each pedestrian.
The tracker maintains a state vector $s_{k}=[x_{k},y_{k},\dot{x}_{k},\dot{y}_{k}]$ encoding position and velocity for each target. At each timestep, Kalman filters first predict target positions using constant velocity dynamics, then the Hungarian algorithm optimally matches these predictions with new detections by minimizing the total Euclidean distance between corresponding points. Unmatched detections initialize new tracks with zero initial velocity, while tracks failing to match consecutively for several frames are terminated to prevent ghost tracking. This pipeline enables robust pedestrian detection and tracking, providing bounding boxes and tracking IDs necessary for subsequent social group detection.
@@
Social Group Detection with Visual Prompting: We leverage a LMM 111specifically gpt-4o-2024-08-06 to identify social groups within crowds by analyzing human behaviors and interactions. Our approach consists of three main steps: image annotation, LMM querying, and social group relationship extraction as shown in Fig. 4. First, given bounding boxes and tracking IDs of pedestrians, We use the mask result from the YOLO-V5, creating an annotated image where each individual is marked with their bounding box, segmentation mask, and unique tracking ID $h_{i}$. These visual marks enhance the LMM’s reasoning. Second, we query the LMM with a structured input consisting of the annotated image and a text prompt. In the text prompt, we provide the LLM with contextual information, including the expected behavior of the LMM, the format of queries and responses, and cues to assist the LLM in accurately identifying social groups. Finally, we process the LMM’s output using a parsing script that extracts the predicted social group relationships. These relationships are then integrated into our tracking system’s data structure, maintaining an up-to-date record of social groups over time. Formally, we can express this process as:
@@
Due to the computational constraints of LMM, we implement a keyframe-based approach to reduce the frequency of LMM queries while maintaining accurate social group tracking. Instead of analyzing every frame, we maintain a keyframe that captures the most informative scene within a 3-second window - specifically, the frame containing the highest number of visible people. Each keyframe stores both the image and corresponding metadata (position and velocity) for all detected individuals. The LMM is only queried when this keyframe is updated, which occurs either when a more representative frame is found or when the current keyframe expires after 3 seconds.
@@ IV-C Socially Aware Planning Module
Algorithm 1 Socially-Aware Mid-level Planning
1:global reference path $R$, cost map $C$, social group $G$
2:CrowdFreePath $F$ or original path $R$
3:$G^{\prime}\leftarrow\text{ConvexHull}(G)$triangleright Compute group boundaries
4:$S\leftarrow\text{EllipsesEnclose}(G^{\prime})$triangleright Generate social spaces
5:$C^{+}\leftarrow\text{UpdateCostmap}(S,C)$triangleright Update cost map
6:if $\text{IsCollided}(R,C^{+})$ then $\triangleright$ Check path collision
7:    $S^{\prime}_{s}\leftarrow\text{GetRobotPose}()$
8:    $S^{\prime}_{g}\leftarrow\text{GetNewGoal}()$
9:    $F\leftarrow\text{BITstarPlanning}(S^{\prime}_{s},S^{\prime}_{g},C^{+})$
10:    return $F$
11:else
12:    return $R$
13:end if
The procedure of the mid-level planning is shown in Algorithm 1. Once the LMM module has produced the social group assignments $G$, we proceed to estimate the spatial extent of each group. Specifically, we compute the convex hull of each group, denoted as $G^{\prime}$, to represent the minimal boundary enclosing all its members. To further approximate the shape and spatial influence of each group, we fit an ellipse to the boundary defined by $G^{\prime}$. The resulting enclosed two-dimensional regions are defined as social spaces $S$, as shown in Fig. 5. These social spaces are then projected onto the environment’s cost map as high-cost regions, producing an updated cost map $C^{+}$ that reflects areas of social sensitivity. If the global reference path $R$ intersects any of the identified social spaces on $C^{+}$, the robot’s current position is designated as a new start point $S^{\prime}_{s}$. A new goal point $S^{\prime}_{g}$ is sampled at a fixed offset from the original reference path such that it lies outside the social space. To compute a socially compliant trajectory under the updated cost constraints, we employ the Batch Informed Trees (BIT*) algorithm [54]. This process yields a revised path, referred to as the CrowdFreePath $F$, which effectively avoids traversing through socially sensitive regions and minimizes potential interactions with social groups. The socially-aware mid-level planning incorporates social group information, enabling dynamic updates to the cost map. This ensures the robot navigates safely while avoiding significant impacts on social groups.
Local Motion Planning: Given the socially-aware path $F$ generated by the mid-level planner, the role of the local motion planner is to generate dynamically feasible trajectories that not only track this path but also adapt in real time to pedestrians. To this end, we develop a local planning framework based on Nonlinear Model Predictive Control with Control Barrier Functions (NMPC-CBF) [55], which explicitly accounts for both the positions and velocities of nearby pedestrians when computing control actions. We define a Control Barrier Function (CBF) as follows:
@@
$\vspace{-0.3cm}\small\min_{\{\mathbf{x}_{k},\mathbf{u}_{k}\}}$ $\quad\|\mathbf{x}_{N}-\mathbf{x}_{\text{goal}}\|_{P_{f}}^{2}+\sum_{k=0}^{N-1}\|\mathbf{u}_{k}\|_{Q_{u}}^{2}$ (4)

### v3

@@ Abstract
With the increasing presence of service robots and autonomous vehicles in human environments, navigation systems need to evolve beyond simple destination reach to incorporate social awareness. This paper introduces GSON, a novel group-based social navigation framework that leverages Large Multimodal Models (LMMs) to enhance robots’ social perception capabilities. Our approach uses visual prompting to enable zero-shot extraction of social relationships among pedestrians and integrates these results with robust pedestrian detection and tracking pipelines to overcome the inherent inference speed limitations of LMMs. The planning system incorporates a mid-level planner that sits between global path planning and local motion planning, effectively preserving both global context and reactive responsiveness while avoiding disruption of the predicted social group. We validate GSON through extensive real-world mobile robot navigation experiments involving complex social scenarios such as queuing, conversations, and photo sessions. Comparative results show that our system significantly outperforms existing navigation approaches in minimizing social perturbations while maintaining comparable performance on traditional navigation metrics.
Index Terms:
Semantic Scene Understanding, AI-Enabled Robotics, Human-Centered Robotics, Vision-Based Navigation
@@ I Introduction
The growth of service robots has driven significant research on autonomous systems capable of navigating human-centered environments [1, 2, 3]. However, a critical gap exists in current navigation systems: while they excel at trajectory prediction and obstacle avoidance [4, 5, 6, 7, 8], they often fail to recognize and respect complex social contexts within crowds, such as photography sessions or queuing behaviors, as illustrated in Fig. 1. In the broader context of social robot navigation [9, 10], the goal is not only for the robot to reach its destination but also to interact appropriately with humans without degrading their experience. In this work, we focus on developing a robot navigation system that can understand and respect contextual social interactions, which we represent as social groups. These social groups represent meaningful patterns of semantic interaction among individuals within the crowd.
Figure 1: Example of a social navigation scenario: Rather than walking between two individuals in a queue, the robot correctly identifies the social context and navigates around the queue to avoid disturbance.
The core issue in building such a socially aware navigation system is how to accurately identify contextual social interactions in dynamic and unpredictable crowds, and exploit these social interactions to guide the planning system. For perception, accurately identifying contextual social interactions in dynamic, unpredictable crowds requires semantic understanding beyond simple proximity or velocity patterns. Traditional methods based on predefined rules or in-domain learning [4, 5, 11, 12, 13, 14, 15, 16] struggle in open-world environments due to data scarcity and high visual variability. Regarding planning, generating socially appropriate motions that respect these interactions demands planning at the right abstraction level—traditional local planners often fall into local minima due to short planning horizons [17, 18, 19, 20], while global path planners cannot adapt quickly to dynamic social interactions. In contrast to modular systems, end-to-end solutions aim to bypass the traditional perception-planning decomposition [21, 22, 23, 24]. However, these approaches often overfit demonstrations and struggle to interpret complex contextual information. More recently, some works have explored the use of large multimodal models (LMMs) to select sampled trajectories [25] or high-level directions[26, 27, 28] to improve social awareness without explicit social interactions modeling. However, these methods are inefficient as they require querying the LMM at every planning step, and the resulting social behaviors are often not interpretable.
@@ II Related Work
Significant research has focused on social group detection in crowded pedestrian environments, which can be approached from both exo-centric and ego-centric perspectives. The exo-centric approach [29] employs external sensors to provide a comprehensive top-down view for crowd analysis. Various methods have been proposed for exo-centric group detection to support motion tracking [30, 31], trajectory prediction[12, 32], group activity recognition[33, 34], and crowd analysis [35, 36, 37]. However, this approach has two key limitations: it relies on external sensors and offline processing, and it often overlooks the rich semantics of crowd interactions by focusing mainly on motion patterns or domain-specific overfitting. For ego-centric group detection, prior approaches to crowd grouping have employed clustering based on motion patterns [38, 13, 14] and distance [39, 40, 14], or learning group information in a specific dataset using graph neural network [15, 16], predicting cost function[41], or leveraging F-formation theories[42]. While these methods excel at robot-centric perception, they still face limitations in capturing semantic context and tend to overfit specific scenarios. In GSON, we leverage the zero-shot visual reasoning capabilities of the Large Multimodal Model (LMM) to enable semantic reasoning of social groups given ego-centric observation.
@@ II-B Social Robot Navigation
Previous studies have proposed various methods to improve collision avoidance in robot navigation [18, 5, 6, 8]. With the increasing number of mobile robots in human environments, there’s a need beyond collision avoidance to improve social awareness. Traditionally, researchers have applied physics-inspired approaches such as the Social Force Model [43] to robot navigation [4, 44]. However, SFM-based methods only implicitly represent social awareness as pedestrian motion and suffer from limited generalization due to parameter selection. Group-based navigation strategies integrate dynamic human grouping into social navigation. Methods that combine group information with model-based planners [38, 40, 14] or reinforcement learning controllers [45, 16] show that representing human groups improves robot safety and social awareness. However, most group-based methods define groups only in terms of position and velocity, ignoring the semantic social interaction between pedestrians. Another line of research uses imitation learning [21, 22, 23, 24] to implicitly learn social awareness from demonstrations. However, due to the data scarcity problem, the learned policies capture only limited behaviors in the dataset. Recent advances in foundation models have integrated large multimodal models to generate high-level actions [27, 28], score trajectories [25] or directions [26], and distill cost into imitation learning [46]. These methods are inefficient, requiring LMM queries at every planning step, with often uninterpretable social behaviors. Our approach uses LMMs only for low-frequency semantic social group detection when new pedestrians appear in observation, handling high-frequency tracking and planning without querying large models, creating a more practical system.
@@ III Problem Formulation
We consider the problem of social robot navigation in a 2D Euclidean space $\mathbb{R}^{2}$. This space is populated by $n$ dynamic pedestrians $H_{i}$ with a specific social structure. We consider the social structure as $m$ social groups $G_{j}$. The robot is initialized with a 2D occupancy map $M$ at the start position $S_{s}$ and a target position $S_{g}$ in the map. Given the 2D lidar point cloud $P^{t}$ and the RGB image observation $I^{t}$, the goal is to enable the robot to safely navigate to the goal position while minimizing the time it spends trespassing on any social group. The robot must infer the social groups of pedestrians from the limited observation and generate an appropriate trajectory $T$. We focus on designing the method to estimate the social group with the aid of the Large Multimodal Model and integrate the estimated result into the navigation system.
@@ IV-B Social Group Estimation with Large Multimodal Models
The core idea of GSON is to model contextual social interactions in dynamic and unpredictable crowds using social groups, which then guide the planning system. In this context, a social group refers to two or more individuals who are interacting, conversing, or moving together with shared intentions or social connections. To identify these social groups, we first build a robust pipeline for pedestrian detection and tracking. After obtaining detection and tracking results of pedestrians, identifying social groups requires commonsense knowledge about social interaction. Therefore, we leverage the multi-modal reasoning capabilities of the LMM and employ visual prompting to identify social groups.
@@
[…]
After this filtering process, we obtain refined pedestrian positions in 2D coordinates $\mathbf{p}_{i}=[x_{i},y_{i}]$, where the $z$ dimension is omitted as we focus on the ground plane for navigation purposes. Finally, we apply the Hungarian algorithm for data association and Kalman filtering for estimating and predicting the position of each pedestrian with a unique tracking ID, enabling robust tracking for subsequent social group detection. Each pedestrian is modeled using a constant velocity motion model, represented by a state vector $\mathbf{s}_{k}=[x_{k},y_{k},\dot{x}_{k},\dot{y}_{k}]^{\top}$, which encodes both position and velocity in 2D space. At each timestep, Kalman filters first predict target positions using linear constant-velocity dynamics, then the Hungarian algorithm optimally matches these predictions with new detections by minimizing the total Euclidean distance between corresponding points. Unmatched detections initialize new tracks with zero initial velocity, while tracks failing to match consecutively for several frames are terminated to prevent ghost tracking.
This pipeline enables robust pedestrian detection and tracking, providing bounding boxes and tracking IDs necessary for subsequent social group detection. In addition to supporting group detection, the tracking results—specifically each pedestrian’s estimated position and velocity—serve as crucial input to the planning module. These motion states help the planner reason about group formations and predict pedestrian movement, enabling proactive, socially aware trajectory generation that avoids disrupting group interactions.
@@
Social Group Detection with Visual Prompting: We leverage a LMM 111specifically gpt-4o-2024-08-06 to identify social groups within crowds by analyzing human behaviors and interactions. Our approach consists of three main steps: image annotation, LMM querying, and social group relationship extraction as shown in Fig. 4. First, given bounding boxes and tracking IDs of pedestrians, We use the mask that results from YOLO-V5, creating an annotated image where each individual is marked with their bounding box, segmentation mask, and unique tracking ID $h_{i}$. These visual marks enhance the LMM’s reasoning. Second, we query the LMM with a structured input consisting of the annotated image and a text prompt. In the text prompt, we provide the LMM with contextual information, including the expected behavior of the LMM, the format of queries and responses, and cues to assist the LMM in accurately identifying social groups. Finally, we process the LMM’s output using a parsing script that extracts the predicted social group relationships. These relationships are then integrated into our tracking system’s data structure, maintaining an up-to-date record of social groups over time. Formally, we can express this process as:
@@
Due to the computational constraints of LMM, we implement a keyframe-based approach to reduce the frequency of LMM queries while maintaining accurate social group tracking. Instead of analyzing every frame, we maintain a keyframe that captures the most informative scene within a 1-second window - specifically, the frame containing the highest number of visible people. Each keyframe stores both the image and corresponding metadata (position and velocity) for all detected individuals. The LMM is only queried when this keyframe is updated, which occurs either when a more representative frame is found or when the current keyframe expires after 1 second.
@@ IV-C Socially Aware Planning Module
Once the LMM module has produced the social group assignments $G$, we proceed to estimate the spatial extent of each group. Specifically, we compute the convex hull of each group, denoted as $G^{\prime}$, to represent the minimal boundary enclosing all its members. To further approximate the shape and spatial influence of each group, we fit an ellipse to the boundary defined by $G^{\prime}$. The resulting enclosed two-dimensional regions are defined as social spaces $S$, as shown in Fig. 5. Once the global reference path $R=\{\mathbf{p}_{0},\dots,\mathbf{p}_{N}\}$ intersects a social space $\mathcal{S}\subset C^{+}$, the robot’s current state is set as a new start $S^{\prime}_{s}$. Let $\mathbf{p}_{k}$ be the last waypoint within $\mathcal{S}$. From $\mathbf{p}_{k}$, we move a fixed arc-length $h$ along $R$ to define a sub-goal $S^{\prime}_{g}$. If the remaining path is shorter than $h$, set $S^{\prime}_{g}=\mathbf{p}_{N}$; otherwise, find the smallest $m$ such that the arc length from $\mathbf{p}_{k}$ to $\mathbf{p}_{m}$ is at least $h$, and interpolate between $\mathbf{p}_{m-1}$ and $\mathbf{p}_{m}$ to locate $S^{\prime}_{g}$. This ensures $S^{\prime}_{g}$ lies on $R$ but outside $\mathcal{S}$, so only the segment $\overline{S^{\prime}_{s}S^{\prime}_{g}}$ requires re-planning.
To compute a socially compliant trajectory under the updated cost constraints, we employ the Batch Informed Trees (BIT*) algorithm [50]. This process yields a revised path, referred to as the CrowdFreePath $F$, which effectively avoids traversing through socially sensitive regions and minimizes potential interactions with social groups. The socially-aware mid-level planning incorporates social group information, enabling dynamic updates to the cost map. This ensures the robot navigates safely while avoiding significant impacts on social groups.
@@
Local Motion Planning: Given the socially aware path $F$ from the mid-level planner and current pedestrian states (positions and velocities), the local planner computes dynamically feasible trajectories that track $F$ and ensure real-time, collision-free navigation by extrapolating pedestrian motion using a constant-velocity model. To this end, we develop a local planning framework based on Nonlinear Model Predictive Control with Control Barrier Functions (NMPC-CBF) [51], which explicitly accounts for both the positions and velocities of nearby pedestrians when computing control actions. We define a Control Barrier Function (CBF) as follows:
@@
$\vskip-8.5359pt\small\min_{\{\mathbf{x}_{k},\mathbf{u}_{k}\}}$ $\quad\|\mathbf{x}_{N}-\mathbf{x}_{\text{goal}}\|_{P_{f}}^{2}+\sum_{k=0}^{N-1}\|\mathbf{u}_{k}\|_{Q_{u}}^{2}$ (4)